Gas Law(What volume does a sample of gas at 60°C and 110kPa occupy if at STP the volume of the gas is 250mL)


1
Expert's answer
2021-05-31T02:01:40-0400

pV = nRT

1) The conditions set by STP is T=273.15 K (0°C or 32°F) and p=105 Pa.

R = 8.3 m3⋅Pa⋅K−1⋅mol−1 = 8.3 L⋅kPa⋅K−1⋅mol−1

The amount of gas under these conditions will be:

n = pV / RT = (105 x 0.00025) / (8.3 x 273.5) = 0.011 mol


2) V = nRT / p = (0.011 x 8.3 x (273 +60)) / 110 = 0.27 L = 270 ml